OpenStudy (anonymous):

domain of sqrt(1-5^t)

OpenStudy (tkhunny):

\[1 - 5^{t} \ge 0\] You tell me why.

OpenStudy (anonymous):

huh

OpenStudy (tkhunny):

That was not a good answer. You WILL have ro do better than that. Find the Domain - Often, it is an exercise in figuring out what doesn't work. Prep Question: What's the square root of -5?

OpenStudy (anonymous):

its nonreal

OpenStudy (tkhunny):

Excellent. What values can we put in a square root and get REAL?

OpenStudy (anonymous):

positive numbers

OpenStudy (tkhunny):

Almost. Let's include zero (0) while we're at it. Now to the problem. We've a square root with an argument. That argument can't be negative. Make sense?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Perfect. Then \[1 - 5^{t} \ge 0\] Still making sense?

OpenStudy (anonymous):

are you saying solve for t

OpenStudy (tkhunny):

Yes. Do you have a plan?

OpenStudy (anonymous):

t is zero

OpenStudy (tkhunny):

t = 0 is part of the solution. It's not an equality. It an INequality.

OpenStudy (anonymous):

all real number

OpenStudy (anonymous):

except t lessthan or equal to 0

OpenStudy (tkhunny):

Now, I think you are guessing a little. Let's use the mathematics. If \[1 - 5^{t} \ge 0\], then \[1\ge 5^{t}\] Thus, we can pick any value of 't' such that \[5^{t}\] does not exceed 1. Still making sense?

OpenStudy (anonymous):

no

OpenStudy (tkhunny):

How did you decide on this? "t lessthan or equal to 0"

OpenStudy (anonymous):

the restrictions in the radicand

OpenStudy (tkhunny):

Well, there you have it.

OpenStudy (anonymous):

I want to put it in interval notation

OpenStudy (tkhunny):

\[(\infty,0]\] That's just a matter of practice.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

how would you write x>ln5 in interval notation

OpenStudy (tkhunny):

\[(ln(5),\infty)\]

OpenStudy (anonymous):

how do you do that

OpenStudy (tkhunny):

Click the [Equation] button and play around with it. Actually, I usually just write the code by hand if it's not too complicated.

OpenStudy (anonymous):

i mean how do you put it in interval notation

OpenStudy (tkhunny):

Well, the interval has endpoints, just list them. Actually, I just lied. The endpoints might be included. 3 < x < 4 is (3,4) 3 < x <= 4 is (3,4] 3 <= x < 4 is [3,4) 3 <= x <= 4 is [3,4] Infinity is NEVER "[" or "]". You can't get there. Always "(" or ")"

OpenStudy (anonymous):

ok but how did you get infinity for x>ln5

OpenStudy (tkhunny):

"Infinity" is a funny word. It means too many things. In this case, it means "there is no limit". Since we've only x > ln(5), with what shall we fill in the other "end point"? ln(5) < x < ??

OpenStudy (anonymous):

ok thanks